OpenStudy (anonymous):

If the complex number x = 3 + bi and |x|^2 = 13, which is a possible value of b? a. 2 b. 4 c. 9 d. 10

OpenStudy (amoodarya):

(3+bi)*(3-bi)=3^2+b^2 =13 so b^2=13-9=4 b=+2,-2 both of them are acceptable